solve for x 2=4+2x

Found 2 solutions by jim_thompson5910, dfrazzetto:
Answer by jim_thompson5910(35256)

2 = 4 + 2x

2 - 4 = 4 + 2x - 4

-2 = 2x

-2/2 = 2x/2

-1 = x

x = -1


So the solution is x = - 1

Answer by dfrazzetto(283)

4+2x = 2
2x = 2-4
2x = -2
x = -1